# GateLens: A Reasoning-Enhanced LLM Agent for Automotive Software Release Analytics

### v1

@@ 8. Conclusions
In real-world deployment serving 60-80 users, GateLens has demonstrated significant practical value through its user-friendly interface and robust query processing capabilities. Users particularly appreciate the flexibility to input, debug, and refine queries easily, marking a substantial advancement in industrial data interaction.
Our implementation insights highlight the advantages of focusing on prompting techniques and in-context learning rather than fine-tuning approaches. This design choice ensures easy integration of newer LLM models without additional training costs while maintaining domain-specific performance (Brown et al., 2020). As we move forward, this modular architecture opens opportunities for incorporating emerging LLM capabilities while preserving the system’s practical utility in safety-critical industrial applications.
Acknowledgements.

### v2

@@ VII Discussion and Conclusions
In real-world deployment serving 60-80 users, GateLens demonstrates significant practical value through its user-friendly interface and robust query processing capabilities, with users particularly appreciating the flexibility to input, debug, and refine queries easily. This marks a substantial advancement in industrial data interaction, successfully handling complex and ambiguous queries while providing practical support for faster decision-making in safety-critical software release processes. GateLens demonstrates significant practical advancements by reducing analysis time by over 80% while maintaining high accuracy in test result interpretation, impact assessment, and release candidate evaluation.
Our implementation insights highlight the advantages of focusing on training-free and single-pass agent systems by foregrounding the perception phase in the code generation pipeline. This modular architecture opens opportunities for incorporating emerging LLM capabilities while preserving the system’s practical utility in safety-critical industrial applications. A phased deployment program is ongoing and shows that multiple stakeholder groups can be supported, though the evolving roles and analytical needs require continued refinement.
Future work will address current limitations by expanding domain applicability beyond automotive software and testing alternative LLM configurations to enhance reliability across other safety-critical industries. This approach demonstrates the potential for LLM-based solutions to transform industrial data analysis workflows while maintaining the reliability standards required for critical applications.
VIII Threats to Validity
The validity of our findings is subject to several potential threats. First, this framework depends on well-defined, static schemas for accurate query decomposition, which fundamentally limits its effectiveness in environments with incomplete or frequently changing schemas. Second, the generalizability of GateLens beyond the automotive software release domain is limited, as the system relies on domain-specific relational modeling and datasets, necessitating further validation in other safety-critical industries. Third, the benchmarks and query scenarios used for evaluation, derived from historical data and real-world queries, may not fully capture the diversity and complexity of potential use cases, which could impact the robustness of the system in broader deployments. Finally, the system’s performance depends on the specific LLM configurations used, such as GPT-4o and Llama 3.1 70B, and their ability to interpret and generate RA expressions. Future work will address these limitations through broader domain testing, expanded evaluation scenarios, and alternative LLM configurations.
Acknowledgments